# SecRepoBench: Benchmarking Code Agents for Secure Code Completion in Real-World Repositories

### v2

@@ 3.3. Failure Modes Analysis
Secure but Incorrect.  While the agent improves a lot in functional correctness, we still observe a few tasks that fail to implement the correct functionality. While this does not give us further information, it shows the effectiveness of our developer-written unit tests, which agrees with previous research (fu2024constrained) that security checks alone are not enough to test the quality of a code completion.
@@ 4. Related Work
Secure Code Generation Benchmarks.  Many benchmarks (chen2021evaluating; liu2024your; austin2021program; hendrycks2measuring; li2022competition; gu2024cruxeval; white2024livebench; jain2024livecodebench; xia2024top; ding2023crosscodeeval; liurepobench; bogomolov2024long; liang2024can) evaluate LLMs on generating correct code. A few benchmarks evaluate how LLMs generate correct code at the repository level (ding2023crosscodeeval; liurepobench; bogomolov2024long; liang2024can), but they do not evaluate the security of generated code. Earlier secure code generation benchmarks only evaluate the security of generated code, such as the Copilot dataset (pearce2022asleep), SecurityEval (siddiq2022securityeval), CodeLMSec (hajipour2024codelmsec), SafeCoder (he2024instruction), and CyberSecEval (bhatt2023purple). Later, researchers developed benchmarks to evaluate both the security and the correctness of generated code (fu2024constrained; peng2025cweval; yang2024seccodeplt; vero2025baxbench), since it is important to evaluate both criteria at the same time. CWEval (peng2025cweval) has demonstrated that it is imprecise to use static analyzer to evaluate the security of generated code. Therefore, recent benchmarks (yang2024seccodeplt; peng2025cweval; vero2025baxbench) opt for using dynamic security test cases. Concurrent work SecCodePLT (yang2024seccodeplt) has C/C++ code completion tasks in real-world repositories. One of the main differences between the C/C++ tasks of SecCodePLT and our work is the method to evaluate the correctness of completed code. SecCodePLT uses mutated fuzzing inputs as the correctness test suites. On the contrary, we use developer-written unit tests to evaluate the correctness of completed code. Compared to SecCodePLT, our method is more realistic.
Techniques to Generate Secure Code.  Researchers have proposed to use prompt engineering (homoliak2024enhancing), prefix tuning (he2023large), instruction tuning (he2024instruction), specialized decoding (fu2024constrained; li2024cosec), specialized prompt optimization (nazzal2024promsec), and vulnerability repair (pearce2023examining; islam2024code; nazzal2024promsec) to make LLMs more likely to generate secure code. In this paper, we are the first to study the secure coding abilities of code agent frameworks in combination with LLMs. We will leave other secure code generation techniques as future work.
@@ 5. Discussions
Generalization of Test Cases.  We reuse the test cases that work on ground truth vulnerable code for evaluating LLM-generated code, including developer-written unit tests and security test cases found by OSS-Fuzz. We assume that the test cases can be generalized to unseen LLM-generated code. This assumption is also made by state-of-the-art benchmarks (yang2024seccodeplt; vero2025baxbench) that contain tasks to generate self-contained programs. Moreover, our procedure to evaluate the security of LLM-generated code follows the workflow of OSS-Fuzz to run known security test cases.

### v3

@@ 3.3. Failure Modes Analysis
Secure but Incorrect.  While the agent improves a lot in functional correctness, we still observe a few tasks that fail to implement the correct functionality. While this does not give us further information, it shows the effectiveness of our developer-written unit tests, which agrees with previous research (Fu et al., 2024) that security checks alone are not enough to test the quality of a code completion.
@@ 4. Related Work
Secure Code Generation Benchmarks.  Many benchmarks (Chen et al., 2021; Liu et al., 2024a; Austin et al., 2021; Hendrycks et al., 2021; Li et al., 2022; Gu et al., 2024; White et al., 2024; Jain et al., 2024; Xia et al., 2024; Ding et al., 2023; Liu et al., 2024b; Bogomolov et al., 2024; Liang et al., 2024) evaluate LLMs on generating correct code. A few benchmarks evaluate how LLMs generate correct code at the repository level (Ding et al., 2023; Liu et al., 2024b; Bogomolov et al., 2024; Liang et al., 2024), but they do not evaluate the security of generated code. Earlier secure code generation benchmarks only evaluate the security of generated code, such as the Copilot dataset (Pearce et al., 2022), SecurityEval (Siddiq and Santos, 2022), CodeLMSec (Hajipour et al., 2024), SafeCoder (He et al., 2024), and CyberSecEval (Bhatt et al., 2023). Later, researchers developed benchmarks to evaluate both the security and the correctness of generated code (Fu et al., 2024; Peng et al., 2025; Nie et al., 2025; Vero et al., 2025), since it is important to evaluate both criteria at the same time. CWEval (Peng et al., 2025) has demonstrated that it is imprecise to use static analyzer to evaluate the security of generated code. Therefore, recent benchmarks (Nie et al., 2025; Peng et al., 2025; Vero et al., 2025) opt for using dynamic security test cases. Concurrent work SecCodePLT (Nie et al., 2025) has C/C++ code completion tasks in real-world repositories. One of the main differences between the C/C++ tasks of SecCodePLT and our work is the method to evaluate the correctness of completed code. SecCodePLT uses mutated fuzzing inputs as the correctness test suites. On the contrary, we use developer-written unit tests to evaluate the correctness of completed code. Compared to SecCodePLT, our method is more realistic.
Techniques to Generate Secure Code.  Researchers have proposed to use prompt engineering (Homoliak et al., 2024), prefix tuning (He and Vechev, 2023), instruction tuning (He et al., 2024), specialized decoding (Fu et al., 2024; Li et al., 2024), specialized prompt optimization (Nazzal et al., 2024), and vulnerability repair (Pearce et al., 2023; Islam and Najafirad, 2024; Nazzal et al., 2024) to make LLMs more likely to generate secure code. In this paper, we are the first to study the secure coding abilities of code agent frameworks in combination with LLMs. We will leave other secure code generation techniques as future work.
@@ 5. Discussions
Generalization of Test Cases.  We reuse the test cases that work on ground truth vulnerable code for evaluating LLM-generated code, including developer-written unit tests and security test cases found by OSS-Fuzz. We assume that the test cases can be generalized to unseen LLM-generated code. This assumption is also made by state-of-the-art benchmarks (Nie et al., 2025; Vero et al., 2025) that contain tasks to generate self-contained programs. Moreover, our procedure to evaluate the security of LLM-generated code follows the workflow of OSS-Fuzz to run known security test cases.